An open-end mutual fund has the following stocks:

Stock Shares Stock Price
A 12,500 $ 85
B 32,000 14
C 3,500 67
D 79,000 11

The fund has 51,000 shares and liabilities of $128,000. Assume the fund is sold with a front-end load of 3.5 percent. What is the offering price of the fund? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

I have answered the question below

Please up vote for the same and thanks!!!

Do reach out in the comments for any queries

Answer:

First step is to calculate the total value of all stocks

=>

= 12500 * 85 + 32000 * 14 + 3500 * 67 + 79000 * 11 = 2614000

= 2614000 - 128,000/51000 = 48.745

= 48.745 * ( 1 + 0.035) = 50.4511